# Adam Exploits $\ell_{\infty}$-geometry of Loss Landscape via Coordinate-wise Adaptivity

### v1

@@ C.1 Proof for convergence rate of blockwise $\mathtt{Adam}$
$H_{b}$ $\geq\sup_{{\bm{x}},{\bm{\Delta}}}\frac{\sqrt{d_{b}}\left\|\nabla_{(b)}L({\bm{x}}+{\bm{\Delta}})-\nabla_{(b)}L({\bm{x}})\right\|_{2}}{\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}}}$
$=\sup_{{\bm{x}},{\bm{\Delta}}}\frac{\sqrt{d_{b}}\left\|\nabla_{(b),:}^{2}L({\bm{x}}){\bm{\Delta}}\right\|_{2}}{\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}}}$
$=\sup_{{\bm{x}},{\bm{\Delta}}}\frac{\sqrt{d_{b}}\left\|\sum_{b^{\prime}=1}^{B}\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}}}$
$=\sup_{{\bm{x}},{\bm{\Delta}},\left\|{\bm{\Delta}}^{\prime}_{(b)}\right\|_{2}\leq 1}\frac{\sqrt{d_{b}}\left\langle{\bm{\Delta}}^{\prime}_{(b)},\sum_{b^{\prime}=1}^{B}\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle}{\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}}}$
$=\sup_{{\bm{x}},\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}\leq\sqrt{d_{b^{\prime}}},\left\|{\bm{\Delta}}^{\prime}_{(b)}\right\|_{2}\leq 1}\sqrt{d_{b}}\left\langle{\bm{\Delta}}^{\prime}_{(b)},\sum_{b^{\prime}=1}^{B}\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle$
$=\sup_{{\bm{x}},{\bm{\Delta}},{\bm{\Delta}}^{\prime}}\sum_{b^{\prime}=1}^{B}\frac{\sqrt{d_{b}}\sqrt{d_{b^{\prime}}}}{\left\|{\bm{\Delta}}^{\prime}_{(b)}\right\|_{2}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}\left\langle{\bm{\Delta}}^{\prime}_{(b)},\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle$
@@
$\frac{H_{b}}{d_{b}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}$ $\geq\frac{\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}}{d_{b}}\sum_{b^{\prime}=1}^{B}\frac{\sqrt{d_{b}}\sqrt{d_{b^{\prime}}}}{\left\|{\bm{\Delta}}_{(b)}\right\|_{2}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}\left\langle{\bm{\Delta}}_{(b)},\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle$
$=\sum_{b^{\prime}=1}^{B}\frac{\sqrt{d_{b^{\prime}}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}}{\sqrt{d_{b}}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}\left\langle{\bm{\Delta}}_{(b)},\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle$
@@
$2\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}$
$=$ $\sum_{b=1}^{B}\frac{H_{b}}{d_{b}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}+\sum_{b^{\prime}=1}^{B}\frac{H_{b^{\prime}}}{d_{b^{\prime}}}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}^{2}$
$\geq$ $\sum_{b=1}^{B}\sum_{b^{\prime}=1}^{B}\frac{\sqrt{d_{b^{\prime}}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}}{\sqrt{d_{b}}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}\left\langle{\bm{\Delta}}_{(b)},\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle+\sum_{b^{\prime}=1}^{B}\sum_{b=1}^{B}\frac{\sqrt{d_{b}}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}}\left\langle{\bm{\Delta}}_{(b^{\prime})},\nabla_{(b^{\prime}),(b)}^{2}L({\bm{x}}){\bm{\Delta}}_{(b)}\right\rangle$
$\geq$ $2\sum_{b=1}^{B}\sum_{b^{\prime}=1}^{B}{\bm{\Delta}}_{(b)}^{\top}\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}=2{\bm{\Delta}}^{\top}\nabla^{2}L({\bm{x}}){\bm{\Delta}}.$
The last inequality comes from mean inequality. ∎
See 3.13
@@
Next we deal with the first order term by approximating it with a deterministic term. Here we need some new notations. Recall that ${\bm{g}}_{t}$ denotes the gradient of mini-batch $L_{t}({\bm{x}}_{t-1})$ at step $t$. And $\mathbb{E}\left[{\bm{g}}_{t}\middle|{\bm{x}}_{t-1}\right]=\nabla L({\bm{x}}_{t-1})$ because $\mathbb{E}{L_{t}}=L$. The full-batch gradient is $\bar{{\bm{g}}}_{t}=\nabla L({\bm{x}}_{t-1})$. Different kinds of second-order momentum are defined in the following way.
$v_{t,b}$ $=\beta_{2}^{t}\left\|{\bm{g}}_{1,(b)}\right\|_{2}^{2}/d_{b}+\left(1-\beta_{2}\right)\sum_{j=0}^{t-1}\beta_{2}^{j}\left(\left\|{\bm{g}}_{t-j,(b)}\right\|_{2}^{2}\right)/d_{b}$
$\tilde{v}_{t,b}$ $=(1-\beta_{2})\left(\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}\right)+\beta_{2}v_{t-1,b}$
See 3.14
@@
This following Lemma C.1 is to control the growth of $v_{T,b}$ so that the right hand side in Lemma 3.13 is indeed $\Theta\left(T+\frac{\log{T}}{1-\beta_{2}}\right)$ instead of $\Theta(T)$ when all the constants are $\text{poly}(T)$.
@@
$\ln\frac{\mathbb{E}\max_{b\in[B]}v_{T,b}+\epsilon}{v_{0}+\epsilon}$ $\leq 2\ln\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}+\max_{b\in[B]}H_{b}^{2}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}$
@@
From the definition of $v_{t,i}$ and 3.4, we have that
@@
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+2(1-\beta_{2}^{t})\sum_{b=1}^{B}\sigma_{b}^{2}+4(1-\beta_{2}^{t})\max_{b\in[B]}\left\|\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}/d_{b}$
$+$ $4(1-\beta_{2})\mathbb{E}\max_{b\in[B]}\sum_{s=1}^{t}\beta_{2}^{t-s}\left\|\nabla_{(b)}L({\bm{x}}_{s-1})-\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}/d_{b}$
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+2(1-\beta_{2}^{t})\sum_{b=1}^{B}\sigma_{b}^{2}+4(1-\beta_{2}^{t})\max_{b\in[B]}\left\|\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}/d_{b}$
$+$ $4(1-\beta_{2})\mathbb{E}\max_{b\in[B]}\sum_{s=1}^{t}\beta_{2}^{t-s}\frac{H_{b}^{2}}{d_{b}^{2}}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{s-1,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}$
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+2\sum_{b=1}^{B}\sigma_{b}^{2}+4\max_{b\in[B]}\left\|\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}/d_{b}+4(1-\beta_{2})(\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}})\mathbb{E}\sum_{s=1}^{t}\beta_{2}^{t-s}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{s-1,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}.$
@@
$\frac{1}{d_{b^{\prime}}}\left\|{\bm{x}}_{t,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}$ $=\frac{\eta^{2}}{d_{b^{\prime}}}\sum_{\Phi(j)=b^{\prime}}\left|\sum_{s=1}^{t}\frac{g_{s,j}}{\sqrt{v_{s,b^{\prime}}+\epsilon}}\right|^{2}$
$\leq\frac{\eta^{2}}{d_{b^{\prime}}}\sum_{\Phi(j)=b^{\prime}}t\sum_{s=1}^{t}\frac{g_{s,j}^{2}}{v_{s,b^{\prime}}+\epsilon}$
$=\eta^{2}t\sum_{s=1}^{t}\frac{\sum_{\Phi(j)=b^{\prime}}g_{s,j}^{2}/d_{b^{\prime}}}{v_{s,b^{\prime}}+\epsilon}$
$\leq\eta^{2}t\left(t+\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{v_{t,b^{\prime}}+\epsilon}{v_{0,b^{\prime}}+\epsilon}\right)$
$\leq\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\max_{b\in[B]}v_{t,b}+\epsilon}{v_{0}+\epsilon}$
The right hand side is independent of specific block $b$. So we can get that
$\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{t,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}\leq\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\max_{b\in[B]}v_{t,b}+\epsilon}{v_{0}+\epsilon}$
and
$\mathbb{E}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{t,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}$
$\leq$ $\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\mathbb{E}\ln\frac{\max_{b\in[B]}v_{t,b}+\epsilon}{v_{0}+\epsilon}$
$\leq$ $\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\mathbb{E}\max_{b\in[B]}v_{t,b}+\epsilon}{v_{0}+\epsilon}$
$\leq$ $\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\epsilon+C+4(1-\beta_{2})\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\mathbb{E}\sum_{s=1}^{t}\beta_{2}^{t-s}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{s-1,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}}{v_{0}+\epsilon}.$
Define $G=\max_{1\leq t\leq T}\mathbb{E}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{t,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}$. There exists $t\leq T$ such that
$G$ $=\mathbb{E}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{t,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}$
$\leq\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\epsilon+C+4(1-\beta_{2})\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\mathbb{E}\sum_{s=1}^{t}\beta_{2}^{t-s}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{x}}_{s-1,(b^{\prime})}-{\bm{x}}_{0,(b^{\prime})}\right\|_{2}^{2}}{d_{b^{\prime}}}}{v_{0}+\epsilon}$
$\leq\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\epsilon+C+4(1-\beta_{2})\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\sum_{s=1}^{t}\beta_{2}^{t-s}G}{v_{0}+\epsilon}$
$\leq\eta^{2}T^{2}+\eta^{2}T\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}G}{v_{0}+\epsilon}.$
We further define $G^{\prime}=\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}G$ and get that
$G^{\prime}$ $\leq\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{1-\beta_{2}}\ln{\frac{G^{\prime}}{v_{0}+\epsilon}}$
$\leq\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}$
$\phantom{\leq\epsilon}+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{1-\beta_{2}}\left(\ln{\frac{G^{\prime}(1-\beta_{2})}{4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\beta_{2}}}+\ln{4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}}\right)$
$\leq\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}+\frac{G^{\prime}}{2}+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{1-\beta_{2}}\ln{4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}}$
$\leq\epsilon+C+4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T^{2}+\frac{G^{\prime}}{2}+(v_{0}+\epsilon)\left(4\max_{b\in[B]}\frac{H_{b}^{2}}{d_{b}^{2}}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}\right)^{2}.$

### v3

@@ Appendix C Proof for Convergence Rate of Blockwise $\mathtt{Adam}$
$H_{b}$ $\geq\sup_{{\bm{x}},{\bm{\Delta}}}\frac{\left\|\nabla_{(b)}L({\bm{x}}+{\bm{\Delta}})-\nabla_{(b)}L({\bm{x}})\right\|_{2}}{\sqrt{d_{b}}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}}}$
$=\sup_{{\bm{x}},{\bm{\Delta}}}\frac{\left\|\nabla_{(b),:}^{2}L({\bm{x}}){\bm{\Delta}}\right\|_{2}}{\sqrt{d_{b}}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}}}$
$=\sup_{{\bm{x}},{\bm{\Delta}}}\frac{\left\|\sum_{b^{\prime}=1}^{B}\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b}}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}}}$
$=\sup_{{\bm{x}},{\bm{\Delta}},\left\|{\bm{\Delta}}^{\prime}_{(b)}\right\|_{2}\leq 1}\frac{\left\langle{\bm{\Delta}}^{\prime}_{(b)},\sum_{b^{\prime}=1}^{B}\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle}{\sqrt{d_{b}}\max_{b^{\prime}\in[B]}\frac{\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}}}$
$=\sup_{{\bm{x}},\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}\leq\sqrt{d_{b^{\prime}}},\left\|{\bm{\Delta}}^{\prime}_{(b)}\right\|_{2}\leq 1}\frac{1}{\sqrt{d_{b}}}\left\langle{\bm{\Delta}}^{\prime}_{(b)},\sum_{b^{\prime}=1}^{B}\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle$
$=\sup_{{\bm{x}},{\bm{\Delta}},{\bm{\Delta}}^{\prime}}\sum_{b^{\prime}=1}^{B}\frac{\sqrt{d_{b^{\prime}}}}{\sqrt{d_{b}}\left\|{\bm{\Delta}}^{\prime}_{(b)}\right\|_{2}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}\left\langle{\bm{\Delta}}^{\prime}_{(b)},\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle.$
@@
$H_{b}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}$ $\geq\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}\sum_{b^{\prime}=1}^{B}\frac{\sqrt{d_{b^{\prime}}}}{\sqrt{d_{b}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}\left|\left\langle{\bm{\Delta}}_{(b)},\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle\right|$
$=\sum_{b^{\prime}=1}^{B}\frac{\sqrt{d_{b^{\prime}}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}}{\sqrt{d_{b}}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}\left|\left\langle{\bm{\Delta}}_{(b)},\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle\right|$
@@
$2\sum_{b=1}^{B}H_{b}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}$
$=$ $\sum_{b=1}^{B}H_{b}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}^{2}+\sum_{b^{\prime}=1}^{B}H_{b^{\prime}}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}^{2}$
$\geq$ $\sum_{b=1}^{B}\sum_{b^{\prime}=1}^{B}\frac{\sqrt{d_{b^{\prime}}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}}{\sqrt{d_{b}}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}\left|\left\langle{\bm{\Delta}}_{(b)},\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right\rangle\right|+\sum_{b^{\prime}=1}^{B}\sum_{b=1}^{B}\frac{\sqrt{d_{b}}\left\|{\bm{\Delta}}_{(b^{\prime})}\right\|_{2}}{\sqrt{d_{b^{\prime}}}\left\|{\bm{\Delta}}_{(b)}\right\|_{2}}\left|\left\langle{\bm{\Delta}}_{(b^{\prime})},\nabla_{(b^{\prime}),(b)}^{2}L({\bm{x}}){\bm{\Delta}}_{(b)}\right\rangle\right|$
$\geq$ $2\sum_{b=1}^{B}\sum_{b^{\prime}=1}^{B}\left|{\bm{\Delta}}_{(b)}^{\top}\nabla_{(b),(b^{\prime})}^{2}L({\bm{x}}){\bm{\Delta}}_{(b^{\prime})}\right|\geq 2\left|{\bm{\Delta}}^{\top}\nabla^{2}L({\bm{x}}){\bm{\Delta}}\right|.$
We will use Lemma 3.14 to better control the growth of the sum of second order term. See 3.14
@@
Next we deal with the first order term by approximating it with a deterministic term. Recall the notation defined in Section 3.4. ${\bm{g}}_{t}$ denotes the gradient of mini-batch $L_{t}({\bm{x}}_{t-1})$ at step $t$. And $\mathbb{E}\left[{\bm{g}}_{t}\middle|{\bm{x}}_{t-1}\right]=\nabla L({\bm{x}}_{t-1})$ because $\mathbb{E}{L_{t}}=L$. The full-batch gradient is $\bar{{\bm{g}}}_{t}=\nabla L({\bm{x}}_{t-1})$. Different kinds of second-order momentum are defined in the following way
$v_{t,b}$ $=\beta_{2}^{t}\left\|{\bm{g}}_{1,(b)}\right\|_{2}^{2}/d_{b}+\left(1-\beta_{2}\right)\sum_{j=0}^{t-1}\beta_{2}^{j}\left(\left\|{\bm{g}}_{t-j,(b)}\right\|_{2}^{2}\right)/d_{b},$
$\tilde{v}_{t,b}$ $=(1-\beta_{2})\left(\left\|\bar{{\bm{g}}}_{t,(b)}\right\|_{2}^{2}/d_{b}+\sigma_{b}^{2}\right)+\beta_{2}v_{t-1,b}.$
See 3.15
@@
This following Lemma C.1 is to control the growth of $v_{T,b}$ so that the right hand side in Lemma 3.14 is indeed $\Theta\left(T+\frac{\log{T}}{1-\beta_{2}}\right)$ instead of $\Theta(\frac{T}{1-\beta_{2}})$ when all the constants are $\text{poly}(T)$.
@@
$\ln\frac{\mathbb{E}\max_{b\in[B]}v_{T,b}+\epsilon}{v_{0}+\epsilon}$ $\leq 2\ln\left(1+\frac{\sum_{b=1}^{B}\sigma_{b}^{2}+\left\|\nabla L({\bm{x}}_{0})\right\|_{\Phi}^{2}+\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T(T+\frac{1}{1-\beta_{2}})}{v_{0}+\epsilon}\right)+\ln{32}$
@@
From the definition of $v_{t,b}$ and 3.10, we have that
@@
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+2(1-\beta_{2}^{t})\sum_{b=1}^{B}\sigma_{b}^{2}+4(1-\beta_{2}^{t})\max_{b\in[B]}\left\|\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}/d_{b}+4(1-\beta_{2})\mathbb{E}\sum_{s=1}^{t}\beta_{2}^{t-s}\left\|\nabla L({\bm{x}}_{s-1})-\nabla L({\bm{x}}_{0})\right\|_{2}^{2}$
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+2(1-\beta_{2}^{t})\sum_{b=1}^{B}\sigma_{b}^{2}+4(1-\beta_{2}^{t})\max_{b\in[B]}\left\|\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}/d_{b}+4(1-\beta_{2})\mathbb{E}\sum_{s=1}^{t}\beta_{2}^{t-s}\sum_{b\in[B]}H_{b}^{2}\left\|{\bm{x}}_{s-1,(b)}-{\bm{x}}_{0,(b)}\right\|_{2}^{2}$
$\leq$ $\beta_{2}^{t}\left\|{\bm{v}}_{0}\right\|_{\infty}+2\sum_{b=1}^{B}\sigma_{b}^{2}+4\max_{b\in[B]}\left\|\nabla_{(b)}L({\bm{x}}_{0})\right\|_{2}^{2}/d_{b}+4(1-\beta_{2})\mathbb{E}\sum_{s=1}^{t}\beta_{2}^{t-s}\sum_{b\in[B]}H_{b}^{2}\left\|{\bm{x}}_{s-1,(b)}-{\bm{x}}_{0,(b)}\right\|_{2}^{2}.$
@@
$\frac{1}{d_{b}}\left\|{\bm{x}}_{t,(b)}-{\bm{x}}_{0,(b)}\right\|_{2}^{2}$ $=\frac{\eta^{2}}{d_{b}}\sum_{\Phi(j)=b}\left|\sum_{s=1}^{t}\frac{g_{s,j}}{\sqrt{v_{s,b}+\epsilon}}\right|^{2}$
$\leq\frac{\eta^{2}}{d_{b}}\sum_{\Phi(j)=b}t\sum_{s=1}^{t}\frac{g_{s,j}^{2}}{v_{s,b}+\epsilon}$
$=\eta^{2}t\sum_{s=1}^{t}\frac{\sum_{\Phi(j)=b}g_{s,j}^{2}/d_{b}}{v_{s,b}+\epsilon}$
$\leq\eta^{2}t\left(t+\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{v_{t,b}+\epsilon}{v_{0,b}+\epsilon}\right)$
$\leq\eta^{2}t^{2}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\ln\frac{\max_{b^{\prime}\in[B]}v_{t,b^{\prime}}+\epsilon}{v_{0}+\epsilon}.$
So we can get that
$\mathbb{E}\sum_{b\in[B]}H_{b}^{2}\left\|{\bm{x}}_{t,(b)}-{\bm{x}}_{0,(b)}\right\|_{2}^{2}$
$\leq$ $\eta^{2}t^{2}\sum_{b\in[B]}H_{b}^{2}d_{b}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\sum_{b\in[B]}H_{b}^{2}d_{b}\mathbb{E}\ln\frac{\max_{b^{\prime}\in[B]}v_{t,b^{\prime}}+\epsilon}{v_{0}+\epsilon}$
$\leq$ $\eta^{2}t^{2}\sum_{b\in[B]}H_{b}^{2}d_{b}+\eta^{2}t\frac{\beta_{2}}{1-\beta_{2}}\sum_{b\in[B]}H_{b}^{2}d_{b}\ln\frac{\mathbb{E}\max_{b^{\prime}\in[B]}v_{t,b^{\prime}}+\epsilon}{v_{0}+\epsilon}.$
Define $G=\max_{1\leq t\leq T}\mathbb{E}\max_{b\in[B]}v_{t,b}+\epsilon$. There exists $t\leq T$ such that
$G$ $=\mathbb{E}\max_{b\in[B]}v_{t,b}+\epsilon$
$\leq\epsilon+C+4(1-\beta_{2})\mathbb{E}\sum_{s=1}^{t-s}\beta_{2}^{t-s}\sum_{b\in[B]}H_{b}^{2}\left\|{\bm{x}}_{s-1,(b)}-{\bm{x}}_{0,(b)}\right\|_{2}^{2}$
$\leq\epsilon+C+4(1-\beta_{2})\sum_{s=1}^{t-s}\beta_{2}^{t-s}\left(\eta^{2}(s-1)^{2}\sum_{b\in[B]}H_{b}^{2}d_{b}+\eta^{2}(s-1)\frac{\beta_{2}}{1-\beta_{2}}\sum_{b\in[B]}H_{b}^{2}d_{b}\ln\frac{G}{v_{0}+\epsilon}\right)$
$\leq\epsilon+C+4\eta^{2}T^{2}\sum_{b\in[B]}H_{b}^{2}d_{b}+4\eta^{2}T\frac{\beta_{2}}{1-\beta_{2}}\sum_{b\in[B]}H_{b}^{2}d_{b}\ln\frac{G}{v_{0}+\epsilon}$
$\leq\epsilon+C+4\eta^{2}T^{2}\sum_{b\in[B]}H_{b}^{2}d_{b}$
$\phantom{\leq\epsilon}+4\eta^{2}T\frac{\beta_{2}}{1-\beta_{2}}\sum_{b\in[B]}H_{b}^{2}d_{b}\left(\ln{\frac{G(1-\beta_{2})}{4\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T\beta_{2}}}+\ln{4\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}}\right)$
$\leq\epsilon+C+4\eta^{2}T^{2}\sum_{b\in[B]}H_{b}^{2}d_{b}+\frac{G}{2}+4\eta^{2}T\frac{\beta_{2}}{1-\beta_{2}}\sum_{b\in[B]}H_{b}^{2}d_{b}\ln{4\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}}$
$\leq\epsilon+C+4\eta^{2}T^{2}\sum_{b\in[B]}H_{b}^{2}d_{b}+\frac{G}{2}+(v_{0}+\epsilon)\left(4\sum_{b\in[B]}H_{b}^{2}d_{b}\eta^{2}T\frac{\beta_{2}}{(1-\beta_{2})(v_{0}+\epsilon)}\right)^{2}.$